# Active Diffusion Subsampling

### v1

@@ 5.2 Application in MRI Acceleration
To assess the real-world practicability of ADS, it was evaluated on the popular fastMRI (Zbontar et al., 2018) 4×4$\times$4 × acceleration benchmark for knee MRIs. In this task, one must reconstruct a fully-sampled knee MRI image given a budget of only 25% of the $\kappa$-space measurements, where each $\kappa$-space measurement is a vertical line of width 1 pixel. We compare with existing MRI acceleration methods focused specifically on learning sampling strategies, namely PG-MRI (Bakker et al., 2020), LOUPE (Bahadir et al., 2020), and SeqMRI (Yin et al., 2021), each of which are detailed in Appendix A.6. We use the same data train / validation / test split and data preprocessing as Yin et al. (2021) for comparability. In particular, the data samples are $\kappa$-space slices cropped and centered at $128\times 128$, with $34,732$ train samples, $1,785$ validation samples, and $1,851$ test samples. We train a DDIM on complex-valued image space samples ${\mathbf{x}}\in\mathbb{C}^{128\times 128}$ obtained by computing the inverse Fourier transform of the $\kappa$-space training samples (see Appendix A.2.2 for further training details). The data space is therefore the complex image space, with $\kappa$-space acting as the measurement space. This yields the measurement model ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})\mathcal{F}({\mathbf{x}})+{\mathbf{n}}_{t}$, where $\mathcal{F}$ is the Discrete Fourier Transform, ${\mathbf{n}}_{t}\sim\mathcal{N}_{\mathcal{C}}(0,\sigma^{2}_{\mathbf{y}})$ is complex Gaussian measurement noise, and ${\mathbf{U}}(A_{t})$ is the subsampling matrix selecting samples at indices $\in A_{t}$. ADS proceeds by running Diffusion Posterior Sampling in the complex image domain with guidance from $\kappa$-space measurements through the measurement model, selecting maximum entropy lines in the $\kappa$-space. We observed on data from the validation set that ADS reconstruction performance increases with the number of reverse diffusion steps, although with diminishing returns for as steps increased. This indicates that in applying ADS, one can choose to increase sample quality at the cost of inference time and compute. To showcase the potential for ADS, we chose a large number of steps, $T=10k$. Further, we choose guidance weight $\zeta=0.85$, and sampling schedule $S$ evenly partitioning $[50,2500]$, and an initial action set $A_{\texttt{init}}=\{63\}$, starting with a single central $\kappa$-space line. Reconstructions are evaluated using the structural similarity index measure (SSIM) (Wang et al., 2004) to compare the absolute values of the fully-sampled target image and reconstructed image. The SSIM uses a window size of $7\times 7$ with $k_{1}=0.01$ and $k_{2}=0.03$ as set be the fastMRI challenge. Table 2 shows the SSIM results on the test set, comparing ADS to recent supervised methods, along with a fixed-mask Diffusion Posterior Sampling using the same inference parameters to serve as a strong unsupervised baseline. The fixed-mask used with DPS measures the 8% of lines at the center of the $\kappa$-space, and random lines elsewhere, as used by Zbontar et al. (2018). It is clear from the results that ADS using performs competitively with supervised baselines, and outperforms the fixed-mask diffusion-based approach. Figure 5 shows two reconstructions created by ADS. See Appendix A.7 for a histogram of all SSIMs over the test set for the diffusion-based approaches.
@@ 6 Discussion
While ADS appears to outperform fixed-mask baselines on both MNIST and fastMRI, it is interesting to note that the relative improvement offered by ADS differs significantly between datasets. For line-based image subsampling on MNIST with $25\%$ samples, ADS achieves a 50% reduction in reconstruction error versus a fixed-mask approach (MAE = 0.019 vs 0.037), whereas with line-based $\kappa$-space subsampling for fastMRI with $25\%$ samples, ADS achieves only a 12% relative improvement (SSIM = 91.26 vs 90.13). We find that the size of this performance gap between fixed and active mask design strategies can be explained by examining the distribution of masks designed by ADS on each task (See Appendix A.4). Indeed, the masks designed for fastMRI are very similar, whereas those designed for MNIST typically differ depending on the sample. When mask designs are similar, then fixed masks will perform similarly to actively designed masks. This is in part a feature of the data distribution – for example, most information in $\kappa$-space is contained in the center. Tasks in which one might expect significantly better performance from ADS are therefore those in which optimal masks will be highly sample-dependent.
In conclusion, we have proposed a method for using diffusion models as active subsampling agents without the need for additional training, using a simple, interpretable action selection policy. We show that this method significantly outperforms fixed-mask baselines on MNIST, and competes with existing supervised approaches in MRI acceleration without tasks-specific training. This method therefore takes a step towards transparent active sensing with automatically generated adaptive strategies, decreasing cost factors such as exposure time, and energy usage.
@@ 7 Limitations & Future Work
While experiments in Section 5 evidence some strengths of ADS against baseline sampling strategies, it is not without limitations. For example, the duration of inference in ADS is dependent on that of the diffusion posterior sampling method. Since low latency is essential in active subsampling, future work could aim to accelerate posterior sampling with diffusion models, leading to accelerated ADS. Another limitation is that the number of measurements taken is upper-bounded by the number of reverse diffusion steps $T$; this limitation could be overcome by extending ADS to generate batch designs (Azimi et al., 2010), containing multiple measurements, from a single posterior estimate. Future work applying ADS in diverse domains would also help to further assess the robustness of the method.
Appendix A Appendix / supplemental material
@@ A.2 Training Details
The methods and models are implemented in the Keras 3.1 (Chollet et al., 2015) library using the Jax backend (Bradbury et al., 2018).
A.2.1 MNIST DDIM Training Details
The model is trained using one GeForce RTX 2080 Ti (NVIDIA, Santa Clara, CA, USA) with 11 GB of VRAM. The training run of 500 epochs took one hours to complete on the specified GPU.
@@ A.2.1 MNIST DDIM Training Details
The DDIM architecture is provided by Keras3 at the following URL: https://keras.io/examples/generative/ddim/. We use widths=[32, 64, 128], block_depth=2, diffusion_steps=30, ema_0.999, learning_rate=0.0001, weight_decay=0.0001, loss="mae".
A.2.2 FastMRI DDIM Training Details
This section provides training details for the training of the DDIM model on FastMRI.
Most of the training details are given in Section 5.2. Additionally, to train the DDIM model, we use the AdamW optimizer (Loshchilov and Hutter, 2017) with an emperically chosen learning rate $1\times 10^{-4}$. The model is trained using one GeForce RTX 2080 Ti (NVIDIA, Santa Clara, CA, USA) with 11 GB of VRAM. The other hyperparameters were not tuned and therefore the model was trained only once. The training run of 305 epochs took 28 hours to complete on the specified GPU.
@@ A.2.2 FastMRI DDIM Training Details
The DDIM architecture is provided by Keras3 at the following URL: https://keras.io/examples/generative/ddim/. We use widths=[32, 64, 96, 128], block_depth=2, diffusion_steps=30, ema_0.999, learning_rate=0.0001, weight_decay=0.0001, loss="mae".
A.3 Validation details
A.3.1 Evaluating FastMRI subsampling stategies
To evaluate the subsampling stategies we utilized a cloud compute cluster with job scheduler. The compute cluster utilizes A100 GPUs (NVIDIA, Santa Clara, CA, USA). This allowed us to parallelize the execution of the diffusion model accross multiple GPUs. For the subsampling stategies some minor hyperparameter tuning was done. To evaluate the complete test set sequentially on one A100 GPU takes 30 hours. other than the hyperparameter tuning there were no other (failed) experiments that used compute.
A.3.2 Evaluating MNIST subsampling stategies
The MNIST experiments used 22 GPU hours on a GeForce RTX 2080 Ti (NVIDIA, Santa Clara, CA, USA) with 11 GB of VRAM. Every diffusion step takes 5 ms. For 1000 diffusion steps, 3 methods, 500 test samples and 11 sampling rates this equals 22 hours. See Figure LABEL:fig:mnist-ads for an example of ADS on MNIST.
A.4 ADS Mask Distributions for MNIST and fastMRI

### v2

@@ 5.3 MRI Acceleration
To assess the real-world practicability of ADS, it was evaluated on the popular fastMRI (Zbontar et al., 2018) 4×4$\times$4 × acceleration benchmark for knee MRIs. In this task, one must reconstruct a fully-sampled knee MRI image given a budget of only 25% of the $\kappa$-space measurements, where each $\kappa$-space measurement is a vertical line of width 1 pixel. We compare with existing MRI acceleration methods focused specifically on learning sampling strategies, namely PG-MRI (Bakker et al., 2020), LOUPE (Bahadir et al., 2020), and SeqMRI (Yin et al., 2021), each of which are detailed in Appendix A.9. We use the same data train / validation / test split and data preprocessing as Yin et al. (2021) for comparability. In particular, the data samples are $\kappa$-space slices cropped and centered at $128\times 128$, with $34,732$ train samples, $1,785$ validation samples, and $1,851$ test samples. We train a diffusion model on complex-valued image space samples ${\mathbf{x}}\in\mathbb{C}^{128\times 128}$ obtained by computing the inverse Fourier transform of the $\kappa$-space training samples (see Appendix A.2.3 for further training details). The data space is therefore the complex image space, with $\kappa$-space acting as the measurement space. This yields the measurement model ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})\mathcal{F}({\mathbf{x}})+{\mathbf{n}}_{t}$, where $\mathcal{F}$ is the Discrete Fourier Transform, ${\mathbf{n}}_{t}\sim\mathcal{N}_{\mathcal{C}}(0,\sigma_{\mathbf{n}})$ is complex Gaussian measurement noise, and ${\mathbf{U}}(A_{t})$ is the subsampling matrix selecting samples at indices $\in A_{t}$. ADS proceeds by running Diffusion Posterior Sampling in the complex image domain with guidance from $\kappa$-space measurements through the measurement model, selecting maximum entropy lines in the $\kappa$-space. We observed on data from the validation set that ADS reconstruction performance increases with the number of reverse diffusion steps, although with diminishing returns as steps increased. This indicates that in applying ADS, one can choose to increase sample quality at the cost of inference time and compute. To showcase the potential for ADS, we chose a large number of steps, $T=10k$. Further, we choose guidance weight $\zeta=0.85$, and sampling schedule $S$ evenly partitioning $[50,2500]$, and an initial action set $A_{\texttt{init}}=\{63\}$, starting with a single central $\kappa$-space line. $N_{p}=16$ and $\sigma_{{\mathbf{y}}}=50$ were used to model the posterior distribution. Reconstructions were evaluated using the structural similarity index measure (SSIM) (Wang et al., 2004) to compare the absolute values of the fully-sampled target image and reconstructed image. The SSIM uses a window size of $4\times 4$ with $k_{1}=0.01$ and $k_{2}=0.03$ as set be the fastMRI challenge. Table 2 shows the SSIM results on the test set, comparing ADS to recent supervised methods, along with a fixed-mask Diffusion Posterior Sampling using the same inference parameters to serve as a strong unsupervised baseline. The fixed-mask used with DPS measures the 8% of lines at the center of the $\kappa$-space, and random lines elsewhere, as used by Zbontar et al. (2018). It is clear from the results that ADS performs competitively with supervised baselines, and outperforms the fixed-mask diffusion-based approach. Figure 6 shows two reconstructions created by ADS. See Appendix A.11 for a histogram of all SSIMs over the test set for the diffusion-based approaches. Finally, we note the inference time for this model, which is an important factor in making active sampling worthwhile. Our model for fastMRI (Table 2) uses $\sim$ 40ms / step with 76 steps per acquisition, leading to 3040ms per acquisition on our NVIDIA GeForce RTX 2080 Ti GPU. A typical acquisition time for a k-space line in MRI is 500ms-2500ms, or higher, depending on the desired quality (Jung & Weigel, 2013). Given modern hardware with increased FLOPs, we believe that this method is already near real-time, even without employing additional tricks to accelerate inference, such as quantization (Shang et al., 2023) or distillation (Salimans & Ho, 2022).
@@
5.4 Focused Ultrasound Scan-Line Subsampling
As a final experiment, we explore focused ultrasound scan-line subsampling (Huijben et al., 2020). In focused ultrasound, a set of focused scan-lines spanning a region of tissue is typically acquired, with the lines being displayed side-by-side to create an image. By acquiring only a small subset of these scan-lines, the power consumption and data transfer rate requirements of the probe can be reduced. Successfully applying scan-line subsampling could find applications in (i) extending battery life in wireless ultrasound probes (Baribeau et al., 2020), (ii) reducing bandwidth requirements for cloud-based ultrasound (Federici et al., 2023), and (iii) increasing frame-rates in 3D ultrasound (Giangrossi et al., 2022). In this experiment, we demonstrate using ADS to identify informative subsets of scan-lines, from which the full image can be reconstructed. To this end, we trained a diffusion model on the samples from the CAMUS (Leclerc et al., 2019) dataset, with the train set consisting of 7448 frames from cardiac ultrasound scans across 500 patients, resized to $128\times 128$ in the polar domain. Each column of pixels in the polar domain is taken to represent a single beamformed scan-line, yielding the model ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t}){\mathbf{x}}$, where ${\mathbf{x}}$ is the fully-sampled target image, ${\mathbf{U}}(A_{t})$ is a mask selecting a set of scan-lines, and ${\mathbf{y}}_{t}$ is the set of scan-line measurements selected by $A_{t}$. Inference takes place in the polar domain, and results are scan-converted for visualization. For ADS inference, the parameters chosen were $T=400$ steps, $S$ evenly partitioning the interval $[0,320]$, $N_{p}=16$, $\sigma_{y}=1$, and $\zeta=10$. We benchmark ADS against DPS using random and data-variance fixed-mask strategies on a test set consisting of frames from 50 unseen patients, finding that ADS significantly outperforms both in terms of reconstruction quality metrics. Figure 7 presents the PSNR results along with a sample reconstruction produced by ADS versus random fixed-mask DPS, with further quantitative results and sample outputs provided in Appendix A.6.
Figure 7: In (a), a comparison of selected measurements and reconstructions generated by ADS versus DPS with a random mask on the CAMUS (Leclerc et al., 2019) dataset. The measurement budget is $16/128=12.5\%$ of the scan-lines. (b) shows the PSNR scores across a held-out test set consisting of 875 frames from 50 unseen patients, with measurement budgets of 16 and 32 scan lines, representing subsampling rates of 12.5% and 25%, respectively.
@@ 6 Discussion
While ADS appears to outperform fixed-mask baselines on MNIST, CelebA, and fastMRI, it is interesting to note that the relative improvement offered by ADS is less pronounced in the case of fastMRI. For line-based image subsampling on MNIST with $25\%$ samples, ADS achieves a 50% reduction in reconstruction error versus a fixed-mask approach (MAE = 0.019 vs 0.037), whereas with line-based $\kappa$-space subsampling for fastMRI with $25\%$ samples, ADS achieves only a 12% relative improvement (SSIM = 91.26 vs 90.13). We find that the size of this performance gap between fixed and active mask design strategies can be explained by examining the distribution of masks designed by ADS on each task (See Appendix A.10). Indeed, the masks designed for fastMRI are very similar, whereas those designed for MNIST typically differ depending on the sample. When mask designs are similar, then fixed masks will perform similarly to actively designed masks. This is in part a feature of the data distribution – for example, most information in $\kappa$-space is contained in the center, at lower frequencies. Tasks in which one might expect significantly better performance from ADS are therefore those in which optimal masks will be highly sample-dependent.
Another interesting trend appears in the results when observing the relative improvement of ADS over fixed-mask strategies as a function of the number of samples taken. One might expect that the relative improvement is highest when the subsampling rate is lowest, monotonically decreasing as more samples are taken. In fact, however, we observe that the largest relative improvement appears not at the lowest subsampling rates, but rather at medium rates, e.g. 10% - 50%. One possible explanation for this is the under-performance of diffusion posterior sampling in scenarios with very few measurements, leading to inaccurate estimates of the posterior distribution and therefore sub-optimal sampling strategies. Further study of the performance of posterior sampling methods in cases with very few measurements could provide an interesting direction for future work.
We also observe that data variance sampling outperforms random sampling on MNIST, but not on CelebA. This somewhat surprising result can be explained by noting that the variance in MNIST appears in a highly informative region, namely around the center of the images where the digits are contained. In contrast, the variance in CelebA appears mostly in the background, where backgrounds may exhibit large deviations in brightness and color. Despite the large variance in the backgrounds, they are typically homogeneous blocks of color that can be inpainted well from sparse samples. Hence, taking most samples in this high-variance background region leading to suboptimal sampling strategies and reconstructions.
A final point of discussion is on choosing the number of particles $N_{p}$ for a particular task. We find in Appendix A.8 that 8-16 particles are sufficient for modelling the posterior in CelebA, with diminishing returns for increasing $N_{p}$. In general, we advise choosing $N_{p}$ such that the modes of the posterior distribution can be sufficiently represented, so that regions of disagreement between modes can be identified by the entropy computation. For example, an MNIST dataset containing only 2 digits will require fewer particles than an MNIST dataset containing 10.
In conclusion, we have proposed a method for using diffusion models as active subsampling agents without the need for additional training, using a simple, interpretable action selection policy. We show that this method significantly outperforms fixed-mask baselines on MNIST, CelebA, and ultrasound scan-line subsampling, and competes with existing supervised approaches in MRI acceleration without tasks-specific training. This method therefore takes a step towards transparent active sensing with automatically generated adaptive strategies, decreasing cost factors such as exposure time, and energy usage.
@@ 7 Limitations & Future Work
While experiments in Section 5 evidence some strengths of ADS against baseline sampling strategies, it is not without limitations. For example, the duration of inference in ADS is dependent on that of the diffusion posterior sampling method. Since low latency is essential in active subsampling, future work could aim to accelerate posterior sampling with diffusion models, leading to accelerated ADS. Another limitation is that the number of measurements taken is upper-bounded by the number of reverse diffusion steps $T$; this limitation could be overcome by extending ADS to generate batch designs (Azimi et al., 2010), containing multiple measurements, from a single posterior estimate. Future work applying ADS in diverse domains would also help to further assess the robustness of the method. Finally, as mentioned in Section 6, we note that ADS does not show significant improvements over baseline sampling strategies under very low sampling rates, e.g. $<5\%$. We hypothesize that this may be due to inaccurate posterior sampling with very few measurements, and believe that this presents an interesting direction for future study.
Appendix A Appendix
@@ A.2 Training Details
The methods and models are implemented in the Keras 3.1 (Chollet et al., 2015) library using the Jax backend (Bradbury et al., 2018). The DDIM architecture is provided by Keras3 at the following URL: https://keras.io/examples/generative/ddim/. Each model was trained using one GeForce RTX 2080 Ti (NVIDIA, Santa Clara, CA, USA) with 11 GB of VRAM.
A.2.1 MNIST
@@
The model was trained for 500 epochs with the following parameters: widths=[32, 64, 128], block_depth=2, diffusion_steps=30, ema_0.999, learning_rate=0.0001, weight_decay=0.0001, loss="mae".
A.2.2 CelebA
@@
The model was trained for 200 epochs with the following parameters: widths=[32, 64, 96, 128], block_depth=2, diffusion_steps=30, ema_0.999, learning_rate=0.0001, weight_decay=0.0001, loss="mae".
A.2.3 FastMRI
The training run of 305 epochs with the following parameters: widths=[32, 64, 96, 128], block_depth=2, diffusion_steps=30, ema_0.999, learning_rate=0.0001, weight_decay=0.0001, loss="mae".
A.3 MNIST metrics
# Pixels (%) Random Data Variance ADS (Ours)
10 (.97%) 0.231 (.002) 0.197 (.002) 0.207 (.002)
25 (2.44%) 0.190 (.002) 0.125 (.002) 0.124 (.002)
50 (4.88%) 0.140 (.002) 0.067 (.001) 0.042 (.001)
100 (9.76%) 0.074 (.001) 0.034 (.001) 0.011 (.000)
250 (24.41%) 0.024 (.000) 0.015 (.000) 0.007 (.000)
500 (48.82%) 0.011 (.000) 0.008 (.000) 0.007 (.000)
# Lines (%) Random Data Variance ADS (Ours)
2 (6.25%) 0.197 (.003) 0.152 (.002) 0.148 (.002)
4 (12.5%) 0.143 (.002) 0.086 (.002) 0.071 (.001)
8 (25%) 0.073 (.001) 0.037 (.001) 0.001 (.000)
16 (50%) 0.022 (.001) 0.013 (.000) 0.000 (.000)
24 (75%) 0.010 (.000) 0.0076 (.000) 0.0074 (.000)
Table 3: Mean and (standard error) for the Mean Absolute Error ($\downarrow$) in reconstruction of MNIST samples, using pixel- and line-based subsampling.
A.4 Comparison with Active Deep Probabilistic Subsampling
In this appendix we provide qualitative and quantitative results comparing ADS to a supervised active sampling approach, Active Deep Probabilistic Subsampling (ADPS) by Van Gorp et al. (2021). ADPS performs active sampling by training a Long Short-Term Memory (LSTM) model to produce logits for each sampling location at each iteration. This is trained jointly with a ‘task model’, which performs some task based on the partial observations. This could be, for example, a reconstruction model. We chose to compare ADS to ADPS on the task of reconstructing MNIST digits to facilitate a qualitative comparison of the masks generated by each method. In order to compare on a reconstruction task, we adapted the MNIST classification model provided in the ADPS codebase [^1] https://github.com/IamHuijben/Deep-Probabilistic-Subsampling/tree/master/ADPS_vanGorp2021/MNIST_Classification for reconstruction by replacing their fully-connected classification layers with a large UNet and using an MSE training objective to reconstruct $32\times 32$ pixel MNIST digits from partial observations containing $81/1024\approx 8\%$ of the pixels. We chose a UNet architecture with widths=[32, 64, 128, 256] and block_depth=2 as the reconstruction model for ADPS, making it similar to the UNet denoising network used by ADS. Both models were trained on the MNIST training set and evaluated on 1000 unseen samples from the validation set.
The results are presented in Figures 8, 9, 10, and 11, indicating that ADS outperforms ADPS in terms of reconstruction error, and generated more adaptive and intuitive masks. One reason for the limited adaptivity in the masks generated by ADPS could be the fact that the reconstruction model is trained jointly with the sampling model, leading to a strategy wherein the reconstruction model can depend on receiving certain inputs and not others. We note, however, that the performance of ADPS could change depending on the choice of reconstruction model, loss function used, and other design choices. In general, the major advantage of ADS over ADPS is that it does not need to be re-trained for different choices of forward model or subsampling rates: all that’s needed is a diffusion prior, and sampling schemes can vary freely at inference time.
(a) ADS
(b) ADPS
Figure 8: t-SNE plots for the masks generated by both (a) ADS and (b) ADPS for 1000 unseen samples from the MNIST validation set. The t-SNE plot for ADS appears to exhibit stronger clustering between different digits, indicating that the generated masks are more digit-dependent. ADPS creates a distinct mask for the digit ‘$1$’, but there is much overlap across other digits.
(a) ADS
(b) ADPS
Figure 9: The mean across all masks for each digit generated by (a) ADS and (b) ADPS. ADS has generated masks that trace the digits, whereas ADPS has generated spread-out masks in the central region of the image where digits may appear.
(a) ADS
(b) ADPS
Figure 10: Random samples from the test set with masks, measurements, and reconstructions generated by (a) ADS and (b) ADPS.
Figure 11: Digit-wise comparison of Mean Absolute Error between targets and reconstructions generated by ADS versus ADPS. The error bars on each bar indicate the standard error of the mean.
A.5 CelebA Further Metrics
Table 4: MAE ($\downarrow$) of the posterior mean on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 19.768 (0.043) 20.625 (0.044) 19.569 (0.060)
100 13.623 (0.028) 14.809 (0.028) 11.731 (0.034)
200 8.909 (0.020) 9.786 (0.020) 6.356 (0.018)
300 6.538 (0.014) 7.507 (0.016) 4.426 (0.012)
500 4.320 (0.009) 5.185 (0.010) 2.825 (0.007)
Table 5: SSIM ($\uparrow$) of the posterior mean on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 0.577 (0.001) 0.557 (0.001) 0.567 (0.001)
100 0.687 (0.001) 0.652 (0.001) 0.697 (0.001)
200 0.791 (0.001) 0.759 (0.000) 0.831 (0.001)
300 0.851 (0.000) 0.820 (0.000) 0.889 (0.000)
500 0.911 (0.000) 0.881 (0.000) 0.941 (0.000)
Table 6: LPIPS ($\downarrow$) (Zhang et al., 2018) of the posterior mean on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 0.306 (0.001) 0.320 (0.000) 0.287 (0.001)
100 0.230 (0.000) 0.248 (0.000) 0.190 (0.001)
200 0.152 (0.000) 0.174 (0.000) 0.103 (0.000)
300 0.111 (0.000) 0.131 (0.000) 0.071 (0.000)
500 0.069 (0.000) 0.086 (0.000) 0.043 (0.000)
Table 7: Mean MAE ($\downarrow$) across posterior samples on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 24.486 (0.048) 25.321 (0.050) 23.895 (0.067)
100 17.436 (0.035) 18.493 (0.034) 14.463 (0.040)
200 11.255 (0.025) 12.337 (0.024) 7.704 (0.021)
300 8.230 (0.018) 9.372 (0.020) 5.225 (0.014)
500 5.389 (0.011) 6.477 (0.013) 3.268 (0.008)
Table 8: Mean PSNR ($\uparrow$) across posterior samples on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 15.988 (0.016) 15.752 (0.015) 16.971 (0.022)
100 18.169 (0.017) 17.791 (0.016) 20.828 (0.023)
200 21.237 (0.019) 20.621 (0.016) 26.314 (0.025)
300 23.617 (0.018) 22.604 (0.018) 30.160 (0.026)
500 26.795 (0.018) 25.274 (0.018) 34.546 (0.026)
Table 9: Mean SSIM ($\uparrow$) across posterior samples on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 0.486 (0.001) 0.471 (0.001) 0.481 (0.001)
100 0.599 (0.001) 0.569 (0.001) 0.623 (0.001)
200 0.725 (0.001) 0.691 (0.001) 0.783 (0.001)
300 0.799 (0.000) 0.764 (0.000) 0.859 (0.001)
500 0.877 (0.000) 0.841 (0.000) 0.925 (0.000)
Table 10: Mean LPIPS ($\downarrow$) (Zhang et al., 2018) across posterior samples on targets from the CelebA test set.
Num Samples Random DPS Data Variance DPS ADS
50 0.321 (0.001) 0.329 (0.001) 0.305 (0.001)
100 0.250 (0.000) 0.260 (0.000) 0.208 (0.001)
200 0.172 (0.000) 0.184 (0.000) 0.112 (0.000)
300 0.127 (0.000) 0.141 (0.000) 0.073 (0.000)
500 0.079 (0.000) 0.095 (0.000) 0.041 (0.000)
A.6 Further Results on Ultrasound Scan-Line Subsampling
Table 11: MAE on test samples from the CAMUS dataset.
Num Samples Random DPS Data Variance DPS ADS
16 12.198 (0.002) 12.431 (0.002) 9.533 (0.001)
32 8.520 (0.001) 8.599 (0.001) 6.701 (0.001)
Table 12: PSNR on test samples from the CAMUS dataset.
Num Samples Random DPS Data Variance DPS ADS
16 22.643 (0.002) 22.480 (0.002) 25.473 (0.001)
32 25.893 (0.001) 25.824 (0.002) 29.130 (0.001)
(a) ADS
(b) Random DPS
Figure 12: Targets, measurements, and reconstructions for 5 random samples from the CAMUS test set by (a) ADS and (b) Random fixed-mask DPS. The Mean Absolute Error for each reconstruction is provided.
A.7 Runtime for CelebA
Figure 13: Runtime of ADS on samples from the CelebA dataset with size $128\times 128$ pixels, using an NVIDIA GeForce RTX 2080 Ti GPU. Note the contribution in runtime from DPS, below the dashed purple line, and the contribution from ADS, above the dashed purple line, which scales linearly with the number of measurements. The execution time presented is the mean execution time over $N=20$ runs, with 1 standard deviation indicated by the error bars.
A.8 Effect of varying the number of particles
Figure 14: Reconstruction accuracy on $N=20$ samples from the the CelebA validation set, as measured by the MAE between the target and posterior mean, using a variety of values for $N_{p}$.
A.9 FastMRI Comparison Methods
A.9.1 PG-MRI
A.9.2 LOUPE
A.9.3 SeqMRI
A.10 ADS Mask Distributions for MNIST and fastMRI